OpenStudy (anonymous):

medal for the dankest answer and explanation : given the parabola with equation y=5x^2-3 find the y value when x=4?

OpenStudy (anonymous):

@im.celibate <3

OpenStudy (anonymous):

@baking4lyf

OpenStudy (anonymous):

This should help, i'm busy.. http://www.math-magic.com/algebra/parabola.htm

OpenStudy (anonymous):

wow that was rude

OpenStudy (anonymous):

@Ogzii

OpenStudy (anonymous):

@William14

OpenStudy (anonymous):

@linda360 hugs and medal ?

OpenStudy (anonymous):

@young.money

OpenStudy (anonymous):

no i don't know this but maybe i can find someone who can

OpenStudy (anonymous):

@bubblegumpop<3

OpenStudy (anonymous):

plug 4 in x then solve the equation

OpenStudy (anonymous):

instead of x, use 4

OpenStudy (anonymous):

divide 4 by 5

OpenStudy (anonymous):

hugs for answer ?

OpenStudy (anonymous):

which is 20, then 20^2-3= 40-3=37 y=37

OpenStudy (anonymous):

yes your now my favorite person in the world

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

*bear hugs*

OpenStudy (anonymous):

aw *hugs back*

OpenStudy (anonymous):

*passes blunt* ?